how do i graph y-4=-2(x-3)
Answer by fractalier(6550)

There are a couple of ways to do this...here is one way...from
y-4=-2(x-3)
let's get it in slope-intercept form, y = mx + b...we get
y - 4 = -2x + 6
y = -2x + 10
Now plot the y-intercept, the point (0, 10).
From there, go down two and to the right one (because the slope is -2/1), and plot another point at (1,8).
Do it again and plot at (2,6).
Connect the points.